You expect to receive the following dividends from a stock:

Year Dividend
1 2
2 4
3 6
4

6

Four years from now, you expect the stock to sell for $79. If the required return is 10%, what is the most you would pay for the stock today? Answer to 2 decimal places, for example $100.12.
Answer #1

LIPUUaru PUPIL foc =SUM(E2:E5) А B 1 Year CashFlow PV@10% Present Value $ 2.00 0.9091 $ 1.82 $ 4.00 0.8264 $ 3.31 $ 6.00 | 0.

fx =SUM(E2:E5) А CashFlow E6 в 1 Year 21 32 43 5 4 6+79 PV@10% =1/1.1 =D2/1.1 =D3/1.1 =D4/1.1 Present Value =C2*D2 =C3*D3 =C4